### 14 Mar 2026 v1

# The Reasoning Bottleneck in Graph-RAG: Structured Prompting and Context Compression for Multi-Hop QA

Abstract
Graph-RAG systems achieve strong multi-hop question answering by indexing documents into knowledge graphs, but strong retrieval does not guarantee strong answers. Evaluating KET-RAG, a leading Graph-RAG system, on three multi-hop QA benchmarks (HotpotQA, MuSiQue, 2WikiMultiHopQA), we find that 77% to 91% of questions have the gold answer in the retrieved context, yet accuracy is only 35% to 78%, and 73% to 84% of errors are reasoning failures. We propose two augmentations: (i) SPARQL chain-of-thought prompting, which decomposes questions into triple-pattern queries aligned with the entity-relationship context, and (ii) graph-walk compression, which compresses the context by $\sim$}60% via knowledge-graph traversal with no LLM calls. SPARQL CoT improves accuracy by +2 to +14 pp; graph-walk compression adds +6 pp on average when paired with structured prompting on smaller models. Surprisingly, we show that, with question-type routing, a fully augmented budget open-weight Llama-8B model matches or exceeds the unaugmented Llama-70B baseline on all three benchmarks at ${\sim}12\times$ lower cost. A replication on LightRAG confirms that our augmentations transfer across Graph-RAG systems.
The Reasoning Bottleneck in Graph-RAG:
Structured Prompting and Context Compression for Multi-Hop QA
Yasaman Zarinkia University of Victoria yasamanz@uvic.ca          Venkatesh Srinivasan Santa Clara University vsrinivasan4@scu.edu          Alex Thomo University of Victoria thomo@uvic.ca
1 Introduction
Multi-hop question answering (QA) requires reasoning over multiple pieces of evidence scattered across documents. Graph-based retrieval-augmented generation (Graph-RAG) addresses this by constructing knowledge graphs from document corpora and using graph structure to retrieve relevant context Edge et al. (2024); Huang et al. (2025).
The retrieval-reasoning gap.
Strong retrieval does not guarantee strong answers. Using KET-RAG Huang et al. (2025), a leading Graph-RAG system, with budget open-weight models (Llama-3.1-8B at \$0.05/M tokens, free 384-dim embeddings, total indexing cost $\sim$}\$1 per dataset), we achieve 77% to 91% context coverage across three benchmarks (Table 2), matching the coverage that KET-RAG reports with commercial models (GPT-4o-mini, OpenAI embeddings). Yet overall accuracy ranges from only 23% to 67% for Llama-3.1-8B and 35% to 78% for Llama-3.3-70B. The culprit is context volume: KET-RAG’s retrieved context typically spans $\sim$}10,000 tokens, containing entity descriptions, relationship descriptions, community report summaries, and raw text chunks. The model must locate the relevant facts and chain them together within this large context, a needle-in-a-haystack challenge that intensifies with the number of reasoning hops required. Decomposing errors into retrieval failures (answer absent from context) and reasoning failures (answer present but model incorrect) reveals that reasoning failures account for 73% to 84% of all errors across datasets. This raises a natural question: can inference-time augmentations to the reasoning stage, requiring no retraining or re-indexing, close this gap?
Reasoning augmentations.
Our augmentations apply to any Graph-RAG system whose retrieved context preserves entity-relationship structure. We focus on KET-RAG and verify that the strategies transfer to LightRAG Guo et al. (2024), another prominent open-source Graph-RAG system. To address this bottleneck, we propose two synergistic strategies that require no additional model training:
  1. 1.
    SPARQL CoT prompting3.1): A structured chain-of-thought prompt that asks the model to decompose the question as a SPARQL query over the retrieved facts, leveraging the inherent entity-relationship structure of the knowledge-graph context. This is a single LLM call with a modified prompt; the only additional cost is slightly longer output.
  2. 2.
    Graph-walk context compression3.2): A post-retrieval filtering step that performs breadth-first traversal from question entities through the knowledge graph, retaining only structurally connected context. This reduces input tokens by $\sim$}60% with no LLM calls, directly reducing the needle-in-a-haystack challenge.
Evaluated on three standard multi-hop QA benchmarks Yang et al. (2018); Trivedi et al. (2022); Ho et al. (2020) across multiple model sizes and configurations (§4), our contributions are:
  1. 1.
    A quantitative error decomposition showing that reasoning, not retrieval, is the dominant bottleneck: 73% to 84% of errors occur when the gold answer is present in the retrieved context.
  2. 2.
    SPARQL CoT consistently improves accuracy (+2 to +14 pp), with the largest gains on harder datasets; an ablation confirms that decomposition drives most of the benefit.
  3. 3.
    A systematic analysis revealing a synergy between graph-walk compression and structured prompting: compression most helps when the model is following a reasoning roadmap, with smaller models gaining +6 pp average.
  4. 4.
    The finding that combining structured prompting, graph-walk compression, and question-type routing enables a budget 8B model to match or exceed the unaugmented 70B baseline on all three benchmarks at ${\sim}12\times$ lower cost.
  5. 5.
    A generality check showing that SPARQL CoT transfers to LightRAG (+7.6 pp on 2WikiMHQA), while graph-walk compression’s effectiveness depends on retrieval-pipeline structure rather than context size alone.
2 Background: KET-RAG
KET-RAG Huang et al. (2025) is a multi-granular indexing framework built on Microsoft’s GraphRAG Edge et al. (2024). It reduces the cost of full-corpus entity extraction by combining three retrieval strategies (K-E-T): (K) a text-keyword bipartite graph built from all chunks without LLM calls; (E) LLM-based entity/relationship extraction on a PageRank-selected subset of core chunks ($\beta{=}0.8$), producing a knowledge graph queried via GraphRAG’s local search; and (T) standard text-chunk retrieval via embedding similarity. A parameter $\theta\in[0,1]$ balances the entity and keyword channels within a total budget of $\lambda$ tokens; we use $\theta{=}0.5$, $\lambda{=}12{,}000$ (defaults of the published code).
KET-RAG achieves state-of-the-art results on multi-hop QA, outperforming all competing systems in its evaluation Lewis et al. (2020); Wang et al. (2024); Edge et al. (2024); Sarmah et al. (2024); Gao et al. (2023); Gutiérrez et al. (2024); Guo et al. (2024). More broadly, KET-RAG is an instance of a growing family of Graph-RAG systems, including LightRAG Guo et al. (2024) and GraphRAG Edge et al. (2024), whose retrieved context preserves explicit entity-relationship structure (entity descriptions, relationship triples, community summaries). This structure creates an opportunity: rather than treating the context as flat text, reasoning augmentations can exploit its entity-relationship layout. We validate this generality on LightRAG in Appendix A.1. We replace KET-RAG’s commercial components (GPT-4o-mini, OpenAI embeddings) with budget open-weight alternatives without loss in coverage (§5.1) and focus on closing the reasoning gap.
3 Methods
3.1 SPARQL Chain-of-Thought Prompting
Figure 1: SPARQL CoT prompting. The model generates a SPARQL query whose triple patterns align with the retrieved context, then traces variable bindings to produce the answer in a single LLM call.
Multi-hop questions require the model to identify and chain multiple facts scattered across a large retrieved context. Standard prompting leaves the model to perform this chaining implicitly, often causing it to miss intermediate hops or abstain when the reasoning path is unclear.
We observe that KET-RAG’s retrieved context is inherently structured: it consists of entity descriptions, relationship descriptions (subject-predicate-object triples), community report summaries, and text chunks. This structure closely mirrors a knowledge graph that can be queried with a formal language. We exploit this alignment by introducing a chain-of-thought (CoT) prompting strategy that uses SPARQL Harris and Seaborne (2013), the W3C standard query language for RDF knowledge graphs, as a reasoning formalism. SPARQL’s triple-pattern syntax (?subject predicate ?object) naturally expresses multi-hop entity chains, making it a good fit for decomposing questions over entity-relationship context. Rather than asking the model to “think step by step” in natural language, we ask it to formulate the question as a structured graph query whose triple patterns map directly onto the entity-relationship pairs in the context (Figure 1).
The prompt template is:
Step 1: Write a simple SPARQL query (max 4 triple patterns, plain English predicates, NO URIs, NO FILTER, NO subqueries). Example:
  SELECT ?answer WHERE {
    ?x name "Paradise Creek" .
    ?x tributaryOf ?y .
    ?y tributaryOf ?answer .
Step 2: Trace the variable bindings for each triple pattern through the context.
Step 3: Write FINAL ANSWER: <your answer>
If the answer is not in the context, write FINAL ANSWER: I don’t know
This is a single LLM call: the model generates the SPARQL query, traces it through the context, and produces the answer in one response. The key advantage over generic sub-question decomposition is structural alignment: the triple patterns (e.g., ?x tributaryOf ?y) correspond directly to the relationship descriptions in the retrieved context, converting an open-ended search over $\sim$}10,000 tokens into a structured template-matching task where each variable binding narrows the search space for the next hop. The only additional cost is the longer output (the SPARQL query and step-by-step trace, typically 100–200 extra tokens).
The final answer is extracted via regex matching on FINAL ANSWER: in the model’s response. If the model determines the answer is not in the context, it may respond with “I don’t know,” which we count as an abstention.
3.2 Graph-Walk Context Compression
While KET-RAG achieves high coverage, much of its $\sim$}10,000-token context is topologically distant from the question’s reasoning chain, increasing the needle-in-a-haystack difficulty.
We propose graph-walk context compression (GW), a post-retrieval step that exploits the knowledge graph’s own topology to isolate a compact reasoning subgraph relevant to the query, with no additional LLM or embedding costs. The algorithm distills the context in four phases:
Figure 2: Graph-walk context compression. BFS from question-anchored seed entities identifies structurally connected entities, assembled into compressed context organized by hop level. No LLM calls required.
Phase 1: Seed entity identification.
Question entities are anchored in the retrieved context via a multi-heuristic string-matching pipeline: (a) exact substring matching (case-insensitive), (b) multi-word matching where all content words of a multi-word entity appear in the question, and (c) partial matching where a significant question word ($\geq$5 characters) appears as part of an entity name; common stop words are excluded.
Phase 2: Topological traversal.
A breadth-first search (BFS) is performed from the seed entities through the knowledge graph, up to a depth of 3 hops. Each reached entity is tagged with its hop distance from the nearest seed, establishing a structural proximity ranking.
Phase 3: Co-occurrence expansion.
The BFS-reached entity set is expanded via a semantic bridge: if a text chunk mentions both a BFS-reached entity and an entity not reached by BFS, the latter is added to the context. This recovers entities that are semantically related but not directly connected in the extracted graph.
Phase 4: Priority-based assembly.
The compressed context is assembled within a token budget $B$ (default B=4, [^0] B{=}4{,}000 tokens), prioritized by structural proximity to the question:
  1. 1.
    Entity and relationship descriptions, ordered by hop distance from seeds (closer = higher priority)
  2. 2.
    Tier-1 text chunks: chunks mentioning BFS-reached entities, ranked by mention count
  3. 3.
    Tier-2 text chunks: chunks sharing question keywords but no BFS-reached entities
  4. 4.
    Community report summaries, if present (lowest priority)
Crucially, the output is reorganized by hop level, creating a natural interface with SPARQL CoT (§3.1): the hop-level sections align with the triple-pattern chain in the generated query, letting the model resolve each variable binding within a localized context region rather than scanning the entire input. If no seed entities are matched, the algorithm falls back to the original uncompressed context.
The entire procedure is pure graph traversal and string matching, requiring no LLM calls or embedding computations. In practice, the BFS reaches 14–21 entities on average, and the priority-based assembly achieves $\sim$}60% compression ($\sim$}10,000 $\to$ $\sim$}4,000 tokens). Figure 2 illustrates the four phases.
4 Experimental Setup
4.1 Datasets
We use three standard multi-hop QA benchmarks:
  • HotpotQA Yang et al. (2018): Bridge and comparison questions over Wikipedia paragraphs. Each question comes with 10 paragraphs (2 gold supporting paragraphs + 8 curated distractors). Bridge questions require following a chain of entities (e.g., “The director of film X was born in what city?”), while comparison questions require attribute comparison (e.g., “Which film was released first, X or Y?”).
  • MuSiQue Trivedi et al. (2022): Multi-hop questions requiring 2–4 reasoning steps, constructed by composing single-hop questions into chains. The deeper reasoning chains (up to 4 hops) are the primary source of difficulty: each additional hop compounds the chance of a reasoning failure. Each question comes with 20 paragraphs (2–4 gold supporting paragraphs + 16–18 distractors drawn from gold paragraphs of related questions, making them harder to distinguish than random distractors). Generally considered the hardest of the three benchmarks.
  • 2WikiMultiHopQA Ho et al. (2020): Questions constructed from Wikidata, requiring inference across two Wikipedia articles with the widest range of reasoning types among the three benchmarks (comparison, inference, compositional, and bridge). Each question comes with 10 paragraphs (2 or 4 gold supporting paragraphs + 6–8 distractors), following HotpotQA’s format. KET-RAG’s own evaluation does not include this dataset; we add it because of its type diversity.
For each dataset, we sample N=500N{=}500 questions using a fixed seed (42) and pool all per-question paragraphs to construct the retrieval corpus.
4.2 Model Stack
Table 1 shows our model configurations. All LLM calls use the Groq inference platform, which provides fast inference for open-weight models at low per-token cost.
Table 1: Model stack. The index is built once with the 8B model and reused across all QA configurations.
Component Model Cost
Embeddings MiniLM-L6-v2 (384d) Free
Index LLM Llama-3.1-8B (Groq) \$0.05/M
QA (budget) Llama-3.1-8B (Groq) \$0.05/M
QA (standard) Llama-3.3-70B (Groq) \$0.59/M
The index is model-independent: the knowledge graph and keyword indices do not depend on the QA model, allowing us to build once and vary the QA model without re-indexing. Compared to KET-RAG’s default stack (GPT-4o-mini + OpenAI embeddings), ours uses a $3\times$ cheaper LLM and free embeddings with $4\times$ lower dimensionality, for a total indexing cost of $\sim$}\$1 per dataset.
At inference time, the 8B QA model is ${\sim}12\times$ cheaper per token than the 70B (\$0.05 vs. \$0.59/M tokens), motivating the question of whether augmentations can close the accuracy gap.
4.3 Configurations
We evaluate four QA configurations combining two prompting strategies with optional graph-walk compression:
  1. 1.
    Baseline: The QA model receives the full KET-RAG-retrieved context ($\sim$}10,000 tokens) and answers directly.
  2. 2.
    Baseline + GW: Same as Baseline but with graph-walk compressed context.
  3. 3.
    SPARQL CoT: The QA model receives the full context with the SPARQL chain-of-thought prompt (§3.1).
  4. 4.
    SPARQL CoT + GW: SPARQL CoT prompting applied to graph-walk compressed context.
Each configuration is run with both the 8B and 70B QA models, yielding 8 runs per dataset and 24 runs total across the three benchmarks, for a total of 12,000 question-answer evaluations. All QA calls use temperature 0.3.
4.4 Evaluation
We report accuracy, token-level F1, and exact match (EM) with SQuAD normalization. Accuracy is determined by a heuristic-first pipeline: normalized string matching (lowercasing, stripping articles/punctuation, substring and last-token comparison), with LLM-judged semantic equivalence (Llama-3.1-8B) as fallback for remaining mismatches (e.g., “NYC” vs. “New York City”). A 102-judgment audit with Claude Opus 4.6 found 96.1% agreement with our pipeline. We also report abstain rate and context coverage (whether the gold answer appears in the retrieved context, separating retrieval from reasoning failures). All comparisons are paired on the same 500 questions and retrieved context for each dataset.
5 Results
5.1 Retrieval Quality: Coverage
With our budget components (free 384-dimensional embeddings, \$0.05/M-token indexing LLM), KET-RAG matches or exceeds its published coverage across all three benchmarks (Table 2), despite $3\times$ cheaper indexing LLM and $4\times$ lower embedding dimensionality.
Table 2: Context coverage (%) with budget components compared to KET-RAG Huang et al. (2025) with GPT-4o-mini + OpenAI embeddings.
Dataset Ours KET Low KET High
HotpotQA 90.8 (454/500) 60.2 82.6
MuSiQue 77.2 (386/500) 77.0 79.6
2WikiMHQA 81.0 (405/500)
5.2 QA Performance
Table 3 presents accuracy, F1, and EM across all 24 configurations.
Table 3: Results (%) on 500 paired questions per dataset. “Covered” = gold answer present in context (count in parentheses). Bold = best per column. GW = graph-walk compression. F1/EM use SQuAD normalization; 8B outputs are LLM-normalized to short answers (§4).
All Questions Covered Questions
Hotpot MuSiQue 2Wiki Hotpot (454) MuSiQue (386) 2Wiki (405)
Method Model Acc F1 EM Acc F1 EM Acc F1 EM Acc F1 EM Acc F1 EM Acc F1 EM
Base 8B 67.0 59.5 49.4 23.6 18.0 12.4 31.4 29.1 24.2 69.6 62.2 51.5 27.5 21.2 15.8 32.1 30.1 24.7
+GW 8B 63.6 55.9 46.0 19.4 14.6 9.0 30.4 28.1 24.8 66.3 58.3 48.0 22.6 17.0 11.4 33.1 30.8 26.9
SPARQL 8B 70.6 60.8 49.4 28.8 22.6 15.4 45.6 36.9 31.0 74.4 64.2 52.6 34.8 27.4 19.7 49.9 40.6 33.6
+GW 8B 76.6 62.2 50.6 30.6 22.9 15.8 55.8 41.1 34.8 79.5 64.9 52.9 35.1 26.4 20.0 59.5 44.5 37.5
Base 70B 78.0 67.9 56.6 35.2 26.9 18.6 48.8 43.0 36.8 81.3 70.9 59.0 40.3 32.5 23.9 54.8 47.9 40.7
+GW 70B 79.2 68.6 56.0 39.8 30.8 20.6 53.6 49.0 44.0 83.3 71.9 58.6 46.2 37.3 26.2 61.0 56.1 50.4
SPARQL 70B 80.2 70.6 57.8 43.8 32.8 23.4 61.0 54.9 48.6 83.9 74.3 61.2 53.2 40.4 29.6 66.9 60.0 52.6
+GW 70B 79.6 70.5 57.8 42.2 34.0 25.8 59.8 51.5 44.0 83.5 74.0 60.8 51.4 42.3 32.5 66.7 57.6 49.6
The headline result is that a budget 8B model, augmented with SPARQL CoT and graph-walk compression, surpasses the unaugmented 70B baseline on 2WikiMHQA (55.8% vs. 48.8%) and closes the gap on MuSiQue and HotpotQA, at ${\sim}12\times$ lower per-token cost. SPARQL CoT alone improves every configuration, with gains up to +14.2 pp on 2WikiMHQA for 8B and +12.2 pp for 70B. When both augmentations are applied, 70B + SPARQL reaches the overall best accuracy (61.0% on 2WikiMHQA, 80.2% on HotpotQA). Difficulty tracks reasoning depth as expected: HotpotQA (2-hop, 67-80%), 2WikiMHQA (mixed, 30-61%), MuSiQue (2-4 hop, 19-44%).
The right half of the table conditions on covered questions (gold answer present in context), separating reasoning failures from retrieval gaps; we discuss this in §5.7. We analyze each factor below.
5.3 Effect of SPARQL CoT Prompting
Table 4 isolates the effect of SPARQL CoT by comparing it to the corresponding baseline (without graph-walk compression).
Table 4: Effect of SPARQL CoT: accuracy $\Delta$ ($p$) and abstain $\Delta$ ($p$) vs. baseline, without GW.
Model Hotpot MuSiQue 2Wiki
Accuracy $\Delta$ ($p$)
8B +3.6+3.6 +5.2+5.2 +14.2+14.2
70B +2.2+2.2 +8.6+8.6 +12.2+12.2
Abstain $\Delta$ ($p$)
8B 6.4-6.4 31.4-31.4 24.4-24.4
70B +0.4+0.4 8.2-8.2 4.2-4.2
The accuracy gains correlate strongly with abstain rate reduction: on MuSiQue, 8B abstentions drop from 52.0% to 20.6% (-31.4 pp), suggesting that expressing relationships as triple patterns forces the model to engage with multi-hop chains rather than declining, converting open-ended reasoning into structured pattern matching.
5.4 CoT Ablation: SPARQL vs. Generic Decomposition
We ablate whether the benefit stems from the SPARQL formalism or decomposition in general, comparing against a generic CoT prompt with natural-language sub-questions on 2WikiMHQA (Table 5).
Table 5: CoT ablation on 2WikiMHQA (without GW).
Method 8B Acc 8B Abs 70B Acc 70B Abs
Baseline 31.4 46.0 48.8 34.8
Generic CoT 52.6 25.2 56.4 36.8
SPARQL CoT 45.6 21.6 61.0 30.6
Routing 58.4 14.2 66.4 26.8
Both CoT methods improve significantly over the baseline, confirming decomposition as the primary driver. The ranking depends on model size: 8B prefers generic CoT (+21.2 pp) over SPARQL (+14.2 pp), while 70B prefers SPARQL (+12.2 pp) over generic (+7.6 pp). SPARQL syntax likely imposes overhead on 8B, while 70B leverages the structural match between triple patterns and KET-RAG’s entity-relationship context. SPARQL CoT also reduces abstain rates more consistently (21.6% vs. 25.2% for 8B; 30.6% vs. 36.8% for 70B). Notably, both CoT methods improve accuracy for the 8B model, contrasting with findings that CoT primarily helps larger models Wei et al. (2022).
A question-type routing policy that directs bridge questions to SPARQL and comparison/inference questions to generic CoT raises 8B accuracy to 58.4%, surpassing the 70B baseline (48.8%). Routing also improves 70B to 66.4%, confirming that question-type-aware CoT selection benefits both model sizes.
5.5 Question-Type Routing
Because 8B’s preferred CoT variant depends on question type (§5.4), we deploy the routing policy across all three datasets with graph-walk compression (Table 6).
Table 6: 8B with routing + GW vs. unaugmented 70B baseline on KET-RAG. Routing classifies each question, applies the preferred CoT, and falls back on abstain.
Accuracy (%) Abstain (%)
Dataset 8B 70B $\Delta$ 8B 70B $\Delta$
HotpotQA 79.8 78.0 +1.8+1.8 1.0 5.6 4.6-4.6
MuSiQue 35.6 35.2 +0.4+0.4 6.2 16.2 10.0-10.0
2WikiMHQA 64.8 48.8 $\mathbf{+16.0}$ 7.2 34.8 $\mathbf{-27.6}$
With routing and graph-walk compression, the 8B model matches or exceeds the unaugmented 70B baseline on all three benchmarks at ${\sim}12\times$ lower cost. The gains are largest on 2WikiMHQA (+16.0 pp), where question-type diversity (bridge, comparison, inference, compositional) gives routing the most leverage. Without routing, SPARQL+GW alone surpasses 70B only on 2WikiMHQA (+7.0 pp) while trailing on HotpotQA (-1.4) and MuSiQue (-4.6); routing closes these remaining gaps. Abstain rates also drop to the lowest of any configuration: 1.0% on HotpotQA, 6.2% on MuSiQue, 7.2% on 2WikiMHQA. The routing policy adds only a cheap classifier call (a few tokens) and at most one retry per question, keeping total inference cost well below a single 70B call.
Implementation.
The router is a single 8B call (max 5 output tokens) that classifies each question as bridge, comparison, or inference (exact prompt in Appendix A.2). Bridge questions go to SPARQL CoT; comparison and inference questions go to generic CoT. If the chosen method abstains, the question is retried with the other method. The full pipeline thus requires at most three 8B calls per question: classify, answer, and optionally retry.
5.6 Effect of Graph-Walk Compression
Graph-walk compression shows a more complex interaction with model capability and prompting strategy (Table 7).
Table 7: Effect of graph-walk compression: accuracy $\Delta$ ($p$) vs. non-GW config.
Method Model Hot. Mus. 2Wi. Avg
Base 8B 3.4-3.4 4.2-4.2 1.0-1.0 2.9-2.9
Base 70B +1.2+1.2 +4.6+4.6 +4.8+4.8 +3.5+3.5
SPARQL 8B +6.0+6.0 +1.8+1.8 +10.2+10.2 +6.0+6.0
SPARQL 70B 0.6-0.6 1.6-1.6 1.2-1.2 1.1-1.1
A consistent pattern emerges: GW helps most for SPARQL + 8B (+6.0 pp avg, +10.2 on 2WikiMHQA) and Base + 70B (+3.5 pp avg), but hurts Base + 8B (-2.9 pp). For SPARQL + 70B, GW incurs only a -1.1 pp average loss while reducing input tokens by $\sim$}60% ($\sim$}10,000 $\to$ $\sim$}4,000), a favorable accuracy-cost trade-off given the 70B model’s higher per-token price. This reveals a complementarity: GW helps when the model has structured guidance (SPARQL + 8B) or sufficient capacity to reason over reduced noise (70B baseline); without either, the small model relies on exhaustive context for pattern matching and compression removes useful signal.
5.7 Error Decomposition
To disentangle retrieval and reasoning failures, we decompose errors by whether the gold answer was present in the retrieved context. Table 8 shows the reasoning share of total errors. The “Covered” columns of Table 3 report accuracy, F1, and EM conditioned on coverage across all configurations.
Table 8: Reasoning share of total errors (%): incorrect answers where the gold answer was present in context.
Config Hotpot MuSiQue 2Wiki
8B Baseline 84 73 80
70B + SPARQL CoT 74 64 69
Reasoning failures dominate: on MuSiQue for example, the answer is present in context for 77.2% of questions (Table 2), yet the 8B baseline answers only 23.6% correctly overall, a gap of over 50 points attributable to reasoning failures. SPARQL CoT with 70B closes roughly half this gap (53.2% on covered questions), confirming that the dominant bottleneck in Graph-RAG QA is reasoning, not retrieval.
On HotpotQA, 70B + SPARQL CoT reaches 83.9% on covered questions, approaching the 90.8% coverage ceiling, suggesting that for 2-hop questions the remaining errors are mostly retrieval gaps.
5.8 Analysis by Question Type
Each benchmark annotates questions with reasoning types. Table 9 breaks down SPARQL CoT gains by type; Table 10 shows GW compression effects by type.
Table 9: SPARQL CoT accuracy improvement ($p$) over baseline, by question type. HotpotQA: bridge (398) and comparison (102). 2WikiMHQA: compositional (205), bridge_comparison (121), comparison (120), inference (54). MuSiQue: by hop depth (264/146/90).
HotpotQA 2WikiMHQA MuSiQue
Model Bri. Comp. Comp. BrCmp. Cmp. Inf. 2h 3h 4h
8B +3.0+3.0 +5.9+5.9 +16.6+16.6 +18.2+18.2 +10.9+10.9 +3.7+3.7 +8.7+8.7 0.6-0.6 +4.5+4.5
70B +2.3+2.3 +2.0+2.0 +4.9+4.9 $\mathbf{+28.1}$ +12.5+12.5 +3.7+3.7 +9.4+9.4 +5.5+5.5 +11.1+11.1
The largest gains appear on chain-following types: bridge_comparison questions see +18.2 pp (8B) and +28.1 pp (70B) on 2WikiMHQA, while inference questions that require implicit reasoning show the smallest gains (+3.7 pp). This aligns with the design of SPARQL CoT: triple-pattern decomposition naturally formalizes entity chains, making it most effective when the question structure matches the formalism. On MuSiQue, the 70B model shows substantial SPARQL CoT benefit at all hop depths (+9.4 at 2-hop, +5.5 at 3-hop, +11.1 at 4-hop), with the largest gain at 4-hop.
Table 10: GW accuracy $\Delta$ ($p$) by question type / hop depth. Counts in parentheses.
Dataset Category SPARQL+8B Base+70B
Hotpot Bridge (398) +4.3+4.3 +1.2+1.2
Comparison (102) +12.8+12.8 +1.0+1.0
2Wiki Composit. (205) +5.8+5.8 1.0-1.0
Br-Comp. (121) +11.6+11.6 +17.4+17.4
Comparison (120) +13.3+13.3 +2.5+2.5
Inference00 (54) +16.7+16.7 +3.7+3.7
MuSiQue 2-hop (264) 1.9-1.9 1.2-1.2
3-hop (146) +6.8+6.8 +11.7+11.7
4-hop0 (90) +4.4+4.4 +10.0+10.0
GW compression similarly benefits deeper hops: on MuSiQue, GW is slightly negative at 2-hop but strongly positive at 3-hop (+11.7 for Base+70B) and 4-hop (+10.0). On 2WikiMHQA, GW helps most on inference (+16.7 pp for SPARQL+8B) and bridge_comparison questions (+17.4 pp for Base+70B), where integrating information across multiple entities benefits most from focused context.
6 Discussion
Reasoning, not retrieval, is the bottleneck.
Interestingly, KET-RAG achieves 77% to 91% coverage even with budget components (§5.1), yet accuracy lags far behind: the coverage-to-accuracy gap exceeds 50 percentage points on MuSiQue (§5.7). As the graph index is model-independent, we can build once with inexpensive models and upgrade the QA component incrementally without re-indexing.
Structured decomposition is the primary lever.
Our key findings show that SPARQL CoT improves every configuration we tested (§3.1), including 8B models, contrasting with findings that chain-of-thought primarily benefits models above $\sim$}100B parameters Wei et al. (2022). The ablation (§5.4) shows the benefit comes from decomposition itself; the SPARQL formalism adds further value for 70B, which can exploit structural alignment with the graph context.
How far can a small model go?
8B + SPARQL CoT + GW surpasses the unaugmented 70B baseline on 2WikiMHQA (55.8% vs. 48.8%) at ${\sim}12\times$ lower per-token cost. On MuSiQue, however, 70B retains a +4.6 pp advantage, suggesting that deeper reasoning chains (3–4 hops) require larger models, while 2-hop questions can be scaffolded effectively even with small models. Question-type routing (§5.5) pushes this further: by steering each question to the appropriate CoT variant and retrying on abstention, the fully augmented 8B surpasses unaugmented 70B on all three benchmarks.
Generality beyond KET-RAG.
To test whether these results are system-specific, we replicate the experiment on LightRAG Guo et al. (2024) across all three benchmarks (Appendix A.1). SPARQL CoT consistently improves LightRAG: +7.6 pp on 2WikiMHQA, +5.4 pp on MuSiQue, and +2.2 pp on HotpotQA, confirming that structured reasoning transfers whenever the context contains entity-relationship structure. Graph-walk compression, however, hurts on LightRAG across all three datasets (-8.6/-6.6/-11.0 pp), despite similar context sizes (median $\sim$}13,000 vs. $\sim$}10,600 tokens). The difference is structural: KET-RAG’s $\beta$-stratified retrieval creates a relevance gradient that graph-walk traversal exploits, whereas LightRAG’s hybrid retrieval returns uniformly relevant material that pruning degrades indiscriminately. This suggests SPARQL CoT is a system-agnostic prompting strategy, while GW compression requires a retrieval-pipeline relevance gradient to be effective.
Deployment recommendations.
The results suggest a simple decision rule. For budget deployments (8B), always enable routing with graph-walk compression: it matches or exceeds 70B at ${\sim}12\times$ lower cost, with negligible overhead (one classifier call plus at most one fallback). For quality-maximizing deployments (70B), use SPARQL CoT without compression; GW’s $\sim$}1 pp accuracy cost rarely justifies the token savings at this scale. GW should be avoided only for Base + 8B, where the small model lacks reasoning guidance to navigate compressed context.
7 Related Work
Graph-RAG systems.
Graph-RAG systems construct knowledge graphs from corpora and use graph structure to guide retrieval. A key distinction is whether the retrieved context preserves entity-relationship (E-R) structure: GraphRAG Edge et al. (2024), KET-RAG Huang et al. (2025), and LightRAG Guo et al. (2024) return entity descriptions, relationship triples, and text chunks, making their context amenable to structured prompting and graph-based compression. HybridRAG Sarmah et al. (2024) also combines KG subgraphs with vector-retrieved passages, though it has only been evaluated on financial earnings-call transcripts.
Other systems use graph structure for retrieval but return flat passages: HippoRAG Gutiérrez et al. (2024, 2025), KGP Wang et al. (2024), Clue-RAG Su et al. (2025), and GraphRAG-V Yu et al. (2025). Xiang et al. (2025) identify context inflation as a key failure mode.
All the above target retrieval quality; we study reasoning over the retrieved E-R context with budget models.
Chain-of-thought prompting.
CoT Wei et al. (2022) and structured variants (Program-of-Thought Chen et al. , decomposed prompting Khot et al. (2023)) improve reasoning by externalizing intermediate steps.
Multi-call methods like IRCoT Trivedi et al. (2023) and Self-Ask Press et al. (2023) interleave reasoning with retrieval; these are orthogonal to our single-call setting where Graph-RAG already provides the full entity-relationship context.
SPARQL has been used for knowledge base QA Gu et al. (2023) but never as a CoT scaffold for Graph-RAG.
Context compression.
LLMLingua Jiang et al. (2023) and RECOMP Xu et al. (2023) compress passages via learned models, incurring inference cost. Graph-RAG context has exploitable topology that enables LLM-free compression via BFS. The interaction between compression and structured prompting has not been previously studied.
8 Conclusion
Once retrieval achieves high coverage, the dominant bottleneck in Graph-RAG QA is reasoning. We propose three inference-time augmentations that target this gap: SPARQL CoT (+2 to +14 pp), graph-walk compression, and question-type routing; none requires retraining or re-indexing. Together, they enable a fully augmented 8B model to match or exceed an unaugmented 70B baseline on all three benchmarks at ${\sim}12\times$ lower cost, and a replication on LightRAG confirms cross-system transfer. Code and data are available at https://github.com/thomouvic/graph-rag-qa-pub.
Limitations
Our study has several limitations. First, we evaluate with a single embedding model and two LLM sizes; a finer-grained sweep across model scales and embedding dimensions would better characterize the degradation curve. Second, our context coverage metric uses substring matching, which may overcount (incidental matches) or undercount (paraphrased answers). Third, we use the default $\theta=0.5$ for KET-RAG’s entity-keyword balance, while tuning $\theta$ per dataset might improve results. Finally, our LLM-judged evaluation, while more flexible than exact match, may introduce evaluation noise; we mitigate this with large sample sizes (N=500N{=}500).
Appendix A Appendix
Table 11: LightRAG generality experiment (%) on 500 paired questions per dataset. “Covered” = gold answer present in LightRAG context (count in parentheses). Bold = best per model group. GW = graph-walk compression. F1/EM use SQuAD normalization; 8B outputs are LLM-normalized to short answers (§4).
All Questions Covered Questions
Hotpot MuSiQue 2Wiki Hotpot (450) MuSiQue (374) 2Wiki (417)
Method Model Acc F1 EM Acc F1 EM Acc F1 EM Acc F1 EM Acc F1 EM Acc F1 EM
Base 8B 62.2 55.1 45.2 26.6 19.3 12.8 32.2 29.0 24.6 65.8 57.8 47.3 34.2 23.7 16.6 34.5 31.2 26.6
+GW 8B 54.4 47.8 38.4 25.4 18.1 11.8 35.2 31.9 27.6 56.0 49.1 39.1 30.7 22.0 15.2 35.7 32.2 27.6
SPARQL 8B 66.8 57.5 46.6 37.4 26.9 18.6 57.4 44.3 36.2 69.6 59.3 47.8 46.3 33.0 24.1 61.9 47.1 38.8
+GW 8B 66.2 55.2 44.4 33.6 23.5 15.8 56.4 44.6 37.8 69.3 57.0 45.6 40.1 28.1 20.3 60.2 46.6 39.8
Base 70B 79.8 68.1 55.6 41.6 32.0 24.2 59.8 49.1 41.8 83.3 70.7 57.8 52.1 40.2 31.8 64.5 52.1 44.8
+GW 70B 68.8 59.9 49.4 35.0 25.7 17.2 51.2 45.8 39.2 71.8 62.2 51.1 44.4 32.5 23.0 54.9 49.0 42.0
SPARQL 70B 82.0 71.6 57.8 47.0 36.5 28.0 67.4 58.1 49.0 85.8 74.2 59.8 58.8 46.6 36.9 72.4 62.7 52.8
+GW 70B 64.8 56.5 45.4 36.6 29.7 23.6 50.2 45.0 37.6 67.3 58.2 46.4 47.1 38.4 31.3 53.7 47.8 40.0
This appendix provides supplementary analysis and experimental details referenced in the main paper. §A.1 presents the full LightRAG generality experiment summarized in §6. §A.2 gives the routing classifier prompt used in §5.5.
A.1 Generality: LightRAG Experiment
To test whether SPARQL CoT generalises beyond KET-RAG, we replicate the full experiment on LightRAG Guo et al. (2024), a prominent open-source Graph-RAG system, across all three benchmarks. We use the same models as our KET-RAG setup (Llama-3.1-8B for indexing, both 8B and 70B for QA) and retrieve hybrid (keyword + vector) context for each of the 500 questions, then run all eight QA configurations (4 methods $\times$ 2 models). LightRAG context coverage is 450/500 on HotpotQA, 374/500 on MuSiQue, and 417/500 on 2WikiMHQA.
Table 11 mirrors Table 3 for direct comparison. SPARQL CoT consistently improves all configurations, with larger gains for 8B, confirming that structured decomposition disproportionately benefits smaller models. The overall gains are smaller than on KET-RAG, likely because LightRAG aggregates entities into high-level descriptions during extraction, making triple-pattern decomposition a less precise match for the retrieved context.
Graph-walk compression, by contrast, hurts on LightRAG across all configurations, the opposite of KET-RAG. Because the two systems retrieve similar-sized contexts (median ${\sim}13{,}000$ vs. ${\sim}10{,}600$ tokens), the difference is not one of context length but of context structure: KET-RAG’s retrieval separates core chunks ($\beta{=}0.8$) from peripheral ones, creating a relevance gradient that graph-walk traversal can exploit. LightRAG’s hybrid retrieval, by contrast, returns material without such stratification, so pruning removes relevant content as readily as noise.
These results support our central claim that SPARQL CoT is a system-agnostic prompting strategy: it improves multi-hop QA whenever the context contains entity-relationship structure, regardless of which Graph-RAG system produced it. The GW finding adds a practical caveat: compression is most beneficial when the retrieval pipeline introduces a clear relevance gradient that graph-walk traversal can exploit, rather than being determined by context size alone.
A.2 Routing Classifier Prompt
The question-type classifier (§5.5) uses the following prompt with the same 8B model (max 5 output tokens):
Classify this question as exactly one of: "bridge", "comparison", "inference".
Definitions:
- bridge: answer requires following an entity chain across facts
- comparison: answer compares two entities/values
- inference: answer requires implicit reasoning not a clean entity chain
Reply with exactly one word: bridge or comparison or inference
Question: {question}
If the model returns a label outside {bridge, comparison, inference}, the system defaults to bridge.